OpenStudy (anonymous):

how do i solve log2(x+4)-log2(x-3)=3

OpenStudy (anonymous):

log2(x+4)-log2(x-3)=log2 2^3 log2 (x+4)/(x-3) = log2 8

OpenStudy (callisto):

\[log_2(x+4)-log_2(x-3)=3\]\[log_2\frac{(x+4)}{(x-3)}=log_22^3\]\[\frac{(x+4)}{(x-3)}=2^3\]\[(x+4)=8(x-3)\]Can you solve it?!

OpenStudy (anonymous):

wait so it is (x+4)= 8x-24 or 5x

OpenStudy (callisto):

x+4 = 8x -24 7x =28 x=?

OpenStudy (anonymous):

x=4

OpenStudy (callisto):

I think so.. but still you can check it again using your calculator !!